OpenStudy (anonymous):

Say you have a right triangle with side a = 21 and side b = 220. What is the length of side c?

OpenStudy (anonymous):

use pythag? c^2 = a^2 + b^2 c being the hypotenuse

OpenStudy (anonymous):

20^2+220^2=c^2 then take sqrt of c^2